# AttentionStore: Cost-effective Attention Reuse across Multi-turn Conversations in Large Language Model Serving

### v1

@@ Abstract
Interacting with humans through multi-turn conversations is a fundamental feature of large language models (LLMs). However, existing LLM serving engines for executing multi-turn conversations are inefficient due to the need to repeatedly compute the key-value (KV) caches of historical tokens, incurring high serving costs. To address the problem, this paper proposes AttentionStore, a new attention mechanism that enables the reuse of KV caches (i.e., attention reuse) across multi-turn conversations, significantly reducing the repetitive computation overheads. AttentionStore maintains a hierarchical KV caching system that leverages cost-effective memory/storage mediums to save KV caches for all requests. To reduce KV cache access overheads from slow mediums, AttentionStore employs layer-wise pre-loading and asynchronous saving schemes to overlap the KV cache access with the GPU computation. To ensure that the KV caches to be accessed are placed in the fastest hierarchy, AttentionStore employs scheduler-aware fetching and eviction schemes to consciously place the KV caches in different layers based on the hints from the inference job scheduler. To avoid the invalidation of the saved KV caches incurred by context window overflow, AttentionStore enables the saved KV caches to remain valid via decoupling the positional encoding and effectively truncating the KV caches. Extensive experimental results demonstrate that AttentionStore significantly decreases the time to the first token (TTFT) by up to 88%, improves the prompt prefilling throughput by 8.2$\times$ for multi-turn conversations, and reduces the end-to-end inference cost by up to 56%. For long sequence inference, AttentionStore reduces the TTFT by up to 95% and improves the prompt prefilling throughput by 22$\times$.
@@ 1 Introduction
Engaging in multi-turn conversations with humans is an essential capability of LLMs [56, 52]. These multi-turn conversations help LLMs comprehend context, user intent, and emotional nuances, enhancing their ability to respond appropriately. Based on the ShareGPT data [65, 41], a widely-used real dataset collected from ChatGPT, 78% of conversations involve multiple turns, as analyzed in Section 2.3.
However, executing multi-turn conversations in current LLM serving engines is highly inefficient, as it requires a large number of repetitive computations, incurring high serving costs. During a single turn of conversation, the LLM engine stores intermediate data, key-value (KV) pairs [4, 37, 22], in the limited high-bandwidth memory (HBM) on GPUs. When that conversation ends and the conversation session becomes inactive, the LLM engine generally discards the KV cache associated with that session, to free up space in the HBM for other active sessions. When the session becomes active again, i.e., the user sends the next message in the conversation, the LLM engine computes the whole KV cache again. This leads to repetitive computation of the same KV cache, wasting valuable GPU computation resources. With the number of conversation turns increases, the repetitive computation overhead linearly increases. Our analysis based on ShareGPT shows that up to 98% of the prefilling cost comes from repetitive computation for the KV cache, as presented in Section 2.3.
@@
We implement the AttentionStore and evaluate it using the real ShareGPT dataset [41]. Extensive experimental results demonstrate that AttentionStore significantly decreases the time to the first token (TTFT) by up to 88% and improves the prompt prefilling throughput by 8.2$\times$ for multi-turn conversations. It also reduces the end-to-end inference cost by up to 56%. For long sequence inference, AttentionStore reduces the TTFT by up to 95% and improves the prompt prefilling throughput by 22$\times$. To summarize, this paper makes the following contributions:
@@ 2.2 Autoregressive Generation
Figure 2: (a) Distribution for conversation turn number in ShareGPT. (b) The session length distribution of ShareGPT.
@@ 2.3 Multi-turn Conversation Inference
Figure 4: Recomputation inefficiencies. (a) The average numbers of historical tokens and new tokens in different turns of ShareGPT. (b) The GPU time for prefilling all tokens and only new input tokens in ShareGPT with LLaMA-70B.
@@
Based on the analysis of ShareGPT [65, 41], a real dataset collected from ChatGPT that includes more than 100K conversations, we observe that 78% of conversations are multi-turn, as shown in Figure LABEL:fig:dist_turnnum. Moreover, 72% of conversations have more than 4K tokens as shown in Figure LABEL:fig:dist_tokennum.
However, executing multi-turn conversations in current LLM serving engines is inefficient due to the repetitive computation of KV caches across multiple conversation turns. As shown in Figure LABEL:fig:conversation, in the conversation turn 1, the LLM serving engine generates the KV cache of $q_{1}$ and $a_{1}$. After finishing turn 1, the LLM serving engine discards the KV cache to reclaim the HBM space. In turn 2, the LLM serving engine re-generates the KV cache of $q_{1}$ and $a_{1}$. In turn 3, the KV cache of $q_{1}$, $a_{1}$, $q_{2}$, and $a_{2}$ is re-generated. As the session expands, the historical tokens keep accumulating and the amount of repetitive computation significantly increases. As shown in Figure LABEL:fig:recompute-tokens, with the increase of the conversation turns, the percentage of historical tokens will be more than 98% in a new conversation. The repetitive computation time occupies 98% of the prefilling time (a.k.a., time to the first token) in the new conversation, as shown in Figure LABEL:fig:recompute-costs.
@@ 2.4 Opportunities and Challenges
4) Unexpected invalidation of the saved KV caches. With the number of conversation turns increasing, the historical tokens can exceed the context window limitation. LLM serving engines generally perform token truncation [33, 16] to reduce the input prompt. The truncation has no impact on previous LLM serving engines since they always recompute the KV cache based on the input prompt following truncation. However, the truncation makes the KV caches saved in the KV caching system invalid, since the position of each token is changed after truncation. Thus it cannot match the old embedded positional encoding in the saved KV cache. Such context window overflow can occur with a high probability. As shown in Figure LABEL:fig:dist_tokennum, 84% and 69% of conversation sessions have a context longer than 2K and 4K, respectively. It means that when using the LLaMA-2 family with 4K context window [48], the context window overflow occurs in 69% of conversation sessions. When using the OPT family with 2K context window [62], the context window overflow occurs in 84% of conversation sessions.
@@ 4.1 Experimental Setup
We implement AttentionStore in Pytorch and Python. The host memory and disks are managed in the form of blocks to improve storage utilization, similar to [22]. Our internal storage allocator allocates and deallocates storage blocks on demand. For the model executor, AttentionStore integrates the implementation of popular LLMs such as LLaMA [47] and Falcon [36] using Pytorch [34] and Transformers [54]. NCCL library [31] is applied for synchronization of the parallel GPU workers. Dedicated CUDA streams are used for moving data between the GPUs and the host memory, overlapping the computation with proactive swapping. Separate IO threads migrate data between the host memory and the disks, overlapping the execution with the KV cache migrations.
Models. The experiments evaluate the open-sourced LLaMA-1 model with 65B [47], LLaMA-2 models [48] with 13B, 70B, and Falcon 40B [36]. All the models in our experiments use FP16 with intermediate activation FP32, aligned with prior systems [60, 51]. Unless specified otherwise, LLaMA-13B operates on two GPUs with 24 batches, while LLaMA-65B, LLaMA-70B, and Falcon-40B run on four GPUs, handling 24 batches each.
@@ 4.2 End-to-end Performance
In the end-to-end experiments, we use 9K conversations from ShareGPT and the average number of turns in these conversations is 3.51. Thus the total number of conversation turns is about 32K. We warm up the KV caching system using the first 10K conversation turns and then evaluate the performance on the following 22K turns.
Cache hit rate. We first present the cache hit rate in AS since other performance metrics are closely related to it. Figure 16 shows the KV cache hit rates in AS for various LLMs. AS exhibits high hit rates around 76%, 56% 87%, and 89% for LLaMa-13B, LLaMA-65B, LLaMA-70B and Falcon-40B, respectively. In contrast, we observe a relatively low hit rate of LLaMA-65B. This discrepancy arises due to the larger storage space required by LLaMA-65B for saving KV caches. Given the same available storage space, AS accommodates fewer sessions for LLaMA-65B, thereby limiting the hit rate. Specifically, LLaMA-65B necessitates 2.5MB of space for each token in the KV cache, LLaMA-13B requires 0.78MB, LLaMA-70B and Falcon-40B require only 0.31MB and 0.12MB of space per token due to using the group query attention with a GQA factor of 8 and 16, respectively.
Time to first token (TTFT). TTFT is an important metric for quality of service in LLM serving [35, 3]. It indicates how quickly users start seeing the output of LLMs after entering their prompt. As shown in Figure 16, AS significantly reduces the TTFT by 84%, 50%, 88% and 88% for LLaMA-13B, LLaMA-65B, LLaMA-70B and Falcon-40B respectively, in comparison to RE. This is because AS eliminates a large amount of repetitive computation for generating the KV caches of historical tokens in the prefilling phase. Upon cache hits, the TTFT of AS only relies on the number of newly input tokens in the new conversation turn.
Prefilling throughput. Prefilling throughput is the metric to evaluate the speed of processing the prompt. Figure 16 shows the measured prefilling throughput. We observe that AS delivers remarkable speedups of 6.3$\times$, 2.0$\times$, 8.1$\times$ and 8.2$\times$ for LLaMA-13B, LLaMA-65B, LLaMA-70B, and Falcon-40B respectively, when compared to RE. The improvement of AS on prefilling throughput comes from the reduced prefilling time. AS only prefills the new input of the new conversation. Moreover, AS can load and reuse the historical KV caches from the KV caching system with layer-wise pre-loading optimization. The historical KV cache loading simultaneously occurs with the prefilling on the new input tokens.
@@
GPU time. Figure 16 shows the end-to-end GPU time to finish all inference jobs in the workload. We observe that AS achieves speedups of 2.7$\times$, 1.6$\times$, 2.2$\times$, and 1.7$\times$ for LLaMA-13B, LLaMA-65B, LLaMA-70B and Falcon-40B respectively, compared to RE. The performance improvements of AS are from two aspects, which are mitigation of recomputing KV caches of the historical tokens, and mitigation of recomputing KV caches after context overflow. Regarding the mitigation of re-prefilling, AS efficiently saves the KV cache in the KV caching system and loads it when necessary for historical tokens. On the other hand, RE discards the KV cache once a job is finished, requiring the redoing of prefilling for every job to reproduce the KV cache. In terms of mitigating the recomputation of KV caches after context flow, RE applies token truncation which invalidates the KV cache for each truncation due to the embedded position encoding in the KV caches. This prompts RE to recompute the KV cache based on the truncated historical tokens. In contrast, AS decouples the position information from the KV caches, allowing direct truncation of the KV cache. This approach avoids the recomputation of KV caches that RE requires.
Inference cost. We evaluate the resource cost based on the on-demand price of AWS EC2 instances [5, 6], i.e., \$5/hour per A100 GPU, \$0.0088/hour/GB for DRAM and \$0.000082/hour/GB for SSD. Figure 18 shows the total costs of different methods for completing the workload. Compared to RE, AS achieves significant cost savings for LLaMA-13B, LLaMA-65B, LLaMA-70B, and Falcon-40B, amounting to 56%, 31%, 49%, and 35%, respectively. These cost savings primarily stem from the reduced GPU time, as AS effectively reduces redundant prefilling for historical tokens and recomputation costs during context overflow, as depicted in Figure 16. AS employs cost-effective storage mediums including host memory and disks to cache the KV caches during inactive conversation sessions. The storage cost from the host memory and disks constitutes 16.4%, 8.9%, 9.0%, and 8.9% of the total cost in AS for LLaMA-13B, LLaMA-65B, LLaMA-70B, and Falcon-40B, respectively.
@@ 4.3.3 Scheduler-aware Fetching and Eviction
As shown in Figure 21(a), for the configuration of 128G/2T that indicates 128GB DRAM and 2TB SSD, AS outperforms LRU and FIFO in the overall cache hit rate by 22% and 18%, respectively. With the increased SSD capacity (128G/10T), AS achieves a remarkable hit rate of 76%, surpassing LRU (31%) and FIFO (48%). AS achieves high overall hit rates because AS is aware of the future KV cache access information to avoid evicting the KV caches that will be used in the future. The higher hit rates are translated to the reduced GPU time as shown in Figure 21(b), where AS achieves speedup up to 1.9$\times$. Analyzing the breakdown of hit rate, for the configuration of 128G/2T, LRU and FIFO only achieve 0.2% and 0.4% (too tiny to display) DRAM hit rates, with the remaining 4.2% and 8.9% being disk hit rates. Even with the overall hit rate increasing to 31% and 48% respectively for the larger capacity of 128G/10T, LRU and FIFO still exhibit limited DRAM hit rates of approximately 0.4% and 0.8% respectively. This is because LRU and FIFO lack awareness of future KV cache information and cannot pre-fetch KV caches from disks to host memory, thereby limiting their ability to improve DRAM hit rates. In contrast, AS achieves a cache hit rate of up to 76%, with over 99.9% of the hits occurring in DRAM due to its scheduler-aware policy.

### v2

@@ Abstract
Interacting with humans through multi-turn conversations is a fundamental feature of large language models (LLMs). However, existing LLM serving engines for executing multi-turn conversations are inefficient due to the need to repeatedly compute the key-value (KV) caches of historical tokens, incurring high serving costs. To address the problem, this paper proposes AttentionStore, a new attention mechanism that enables the reuse of KV caches (i.e., attention reuse) across multi-turn conversations, significantly reducing the repetitive computation overheads. AttentionStore maintains a hierarchical KV caching system that leverages cost-effective memory/storage mediums to save KV caches for all requests. To reduce KV cache access overheads from slow mediums, AttentionStore employs layer-wise pre-loading and asynchronous saving schemes to overlap the KV cache access with the GPU computation. To ensure that the KV caches to be accessed are placed in the fastest hierarchy, AttentionStore employs scheduler-aware fetching and eviction schemes to consciously place the KV caches in different layers based on the hints from the inference job scheduler. To avoid the invalidation of the saved KV caches incurred by context window overflow, AttentionStore enables the saved KV caches to remain valid via decoupling the positional encoding and effectively truncating the KV caches. Extensive experimental results demonstrate that AttentionStore significantly decreases the time to the first token (TTFT) by up to 87%, improves the prompt prefilling throughput by 7.8$\times$ for multi-turn conversations, and reduces the end-to-end inference cost by up to 70%. For long sequence inference, AttentionStore reduces the TTFT by up to 95% and improves the prompt prefilling throughput by 22$\times$.
@@ 1 Introduction
Engaging in multi-turn conversations with humans is an essential capability of LLMs [57, 53]. These multi-turn conversations help LLMs comprehend context, user intent, and emotional nuances, enhancing their ability to respond appropriately. Based on the ShareGPT data [37], a widely-used real dataset collected from ChatGPT, 73% of conversations involve multiple turns, as analyzed in Section 2.3.
However, executing multi-turn conversations in current LLM serving engines is highly inefficient, as it requires a large number of repetitive computations, incurring high serving costs. During a single turn of conversation, the LLM engine stores intermediate data, key-value (KV) pairs [4, 38, 22], in the limited high-bandwidth memory (HBM) on GPUs. When that conversation ends and the conversation session becomes inactive, the LLM engine generally discards the KV cache associated with that session, to free up space in the HBM for other active sessions. When the session becomes active again, i.e., the user sends the next message in the conversation, the LLM engine computes the whole KV cache again. This leads to repetitive computation of the same KV cache, wasting valuable GPU computation resources. With the number of conversation turns increases, the repetitive computation overhead linearly increases. Our analysis based on ShareGPT shows that up to 99% of the prefilling cost comes from repetitive computation for the KV cache, as presented in Section 2.3.
@@
We implement the AttentionStore and evaluate it using the real ShareGPT dataset [37]. Extensive experimental results demonstrate that AttentionStore significantly decreases the time to the first token (TTFT) by up to 87% and improves the prompt prefilling throughput by 7.8$\times$ for multi-turn conversations. It also reduces the end-to-end inference cost by up to 70%. For long sequence inference, AttentionStore reduces the TTFT by up to 95% and improves the prompt prefilling throughput by 22$\times$. To summarize, this paper makes the following contributions:
@@ 2.2 Autoregressive Generation
Figure 2: (a) Distribution for conversation turn number in ShareGPT [37]. (b) The session length distribution of ShareGPT. For better display effect, the statistics exclude conversations with over 40 turns or sessions that exceed a length of 32K.
@@ 2.3 Multi-turn Conversation Inference
Figure 4: Recomputation inefficiencies. (a) The average numbers of historical tokens and new tokens in different turns of ShareGPT [37]. (b) The GPU time for prefilling all tokens and only new input tokens in ShareGPT with Mistral-7B [20] on 1 A100 GPU.
@@
Based on the analysis of ShareGPT [37, 42], a real dataset collected from ChatGPT that includes more than 90K conversations, we observe that 73% of conversations are multi-turn, as shown in Figure LABEL:fig:dist_turnnum. Moreover, 30% of conversations have more than 4K tokens as shown in Figure LABEL:fig:dist_tokennum.
However, executing multi-turn conversations in current LLM serving engines is inefficient due to the repetitive computation of KV caches across multiple conversation turns. As shown in Figure LABEL:fig:conversation, in the conversation turn 1, the LLM serving engine generates the KV cache of $q_{1}$ and $a_{1}$. After finishing turn 1, the LLM serving engine discards the KV cache to reclaim the HBM space. In turn 2, the LLM serving engine re-generates the KV cache of $q_{1}$ and $a_{1}$. In turn 3, the KV cache of $q_{1}$, $a_{1}$, $q_{2}$, and $a_{2}$ is re-generated. As the session expands, the historical tokens keep accumulating and the amount of repetitive computation significantly increases. As shown in Figure LABEL:fig:recompute-tokens, with the increase of the conversation turns, the percentage of historical tokens will be more than 99% in a new conversation. The repetitive computation time occupies 99% of the prefilling time (a.k.a., time to the first token) in the new conversation, as shown in Figure LABEL:fig:recompute-costs.
@@ 2.4 Opportunities and Challenges
4) Unexpected invalidation of the saved KV caches. With the number of conversation turns increasing, the historical tokens can exceed the context window limitation. LLM serving engines generally perform token truncation [33, 16] to reduce the input prompt. The truncation has no impact on previous LLM serving engines since they always recompute the KV cache based on the input prompt following truncation. However, the truncation makes the KV caches saved in the KV caching system invalid, since the position of each token is changed after truncation. Thus it cannot match the old embedded positional encoding in the saved KV cache. Such context window overflow can occur with a high probability. As shown in Figure LABEL:fig:dist_tokennum, 47% and 30% of conversation sessions have a context longer than 2K and 4K, respectively. It means that when using the LLaMA-2 family with 4K context window [49], the context window overflow occurs in 30% of conversation sessions. When using the OPT family with 2K context window [63], the context window overflow occurs in 47% of conversation sessions.
@@ 4.1 Experimental Setup
We implement AttentionStore in Pytorch and Python. The host memory and disks are managed in the form of blocks to improve storage utilization, similar to [22]. Our internal storage allocator allocates and deallocates storage blocks on demand. For the model executor, AttentionStore integrates the implementation of popular LLMs such as LLaMA [48] and Falcon [36] using Pytorch [34] and Transformers [55]. NCCL library [31] is applied for synchronization of the parallel GPU workers. Dedicated CUDA streams are used for moving data between the GPUs and the host memory, overlapping the computation with proactive swapping. Separate IO threads migrate data between the host memory and the disks, overlapping the execution with the KV cache migrations. Continuous batching [61] is enabled through experiments.
Models. The experiments evaluate the open-sourced LLaMA-1 model with 65B [48], LLaMA-2 models [49] with 13B, 70B, and Falcon 40B [36]. The intermediate activation uses FP16, aligned with prior systems [61, 52]. We also implement Mistral-7B [20] with a 32K context window. Unless specified otherwise, LLaMA-13B operates on two GPUs with 24 batches, while LLaMA-65B, LLaMA-70B, and Falcon-40B run on four GPUs, handling 24 batches each.
@@ 4.2 End-to-end Performance
In the end-to-end experiments, we use 9K conversations from ShareGPT [37] and the average number of turns in these conversations is 5.75. Thus the total number of conversation turns is about 52K. We warm up the KV caching system using the first 10K conversation turns and then evaluate the performance on the following 42K turns.
Cache hit rate. We first present the cache hit rate in AS since other performance metrics are closely related to it. Figure 16 shows the KV cache hit rates in AS for various LLMs. AS exhibits high hit rates around 86%, 71% 89%, and 90% for LLaMa-13B, LLaMA-65B, LLaMA-70B and Falcon-40B, respectively. In contrast, we observe a relatively low hit rate of LLaMA-65B. This discrepancy arises due to the larger storage space required by LLaMA-65B for saving KV caches. Given the same available storage space, AS accommodates fewer sessions for LLaMA-65B, thereby limiting the hit rate. Specifically, LLaMA-65B necessitates 2.5MB of space for each token in the KV cache, LLaMA-13B requires 0.78MB, LLaMA-70B and Falcon-40B require only 0.31MB and 0.12MB of space per token due to using the group query attention with a GQA factor of 8 and 16, respectively.
Time to first token (TTFT). TTFT is an important metric for quality of service in LLM serving [35, 3]. It indicates how quickly users start seeing the output of LLMs after entering their prompt. As shown in Figure 16, AS significantly reduces the TTFT by 85%, 61%, 87% and 86% for LLaMA-13B, LLaMA-65B, LLaMA-70B and Falcon-40B respectively, in comparison to RE. This is because AS eliminates a large amount of repetitive computation for generating the KV caches of historical tokens in the prefilling phase. Upon cache hits, the TTFT of AS only relies on the number of newly input tokens in the new conversation turn.
Prefilling throughput. Prefilling throughput is the metric to evaluate the speed of processing the prompt. Figure 16 shows the measured prefilling throughput. We observe that AS delivers remarkable speedups of 6.8$\times$, 2.6$\times$, 7.8$\times$ and 7.2$\times$ for LLaMA-13B, LLaMA-65B, LLaMA-70B, and Falcon-40B respectively, when compared to RE. The improvement of AS on prefilling throughput comes from the reduced prefilling time. AS only prefills the new input of the new conversation. Moreover, AS can load and reuse the historical KV caches from the KV caching system with layer-wise pre-loading optimization. The historical KV cache loading simultaneously occurs with the prefilling on the new input tokens.
@@
GPU time. Figure 16 shows the end-to-end GPU time to finish all inference jobs in the workload. We observe that AS achieves speedups of 4.0$\times$, 1.9$\times$, 3.3$\times$, and 3.4$\times$ for LLaMA-13B, LLaMA-65B, LLaMA-70B and Falcon-40B respectively, compared to RE. The performance improvements of AS are from two aspects, which are mitigation of recomputing KV caches of the historical tokens, and the mitigation of recomputing KV caches after context overflow. Regarding the mitigation of re-prefilling, AS efficiently saves the KV cache in the KV caching system and loads it when necessary for historical tokens. On the other hand, RE discards the KV cache once a job is finished, requiring the redoing of prefilling for every job to reproduce the KV cache. In terms of mitigating the recomputation of KV caches after context flow, RE applies token truncation which invalidates the KV cache for each truncation due to the embedded position encoding in the KV caches. This prompts RE to recompute the KV cache based on the truncated historical tokens. In contrast, AS decouples the position information from the KV caches, allowing direct truncation of the KV cache. This approach avoids the recomputation of KV caches that RE requires. Note that AttentionStore also not only aids in minimizing the prefilling time but also helps reduce the decoding time [61]. Specifically, under continuous batching, each newly arrived job must complete prefilling before it can join other decoding jobs. This process blocks the execution of decoding jobs, resulting in prolonged decoding time. However, AttentionStore mitigates this issue by minimizing prefilling time for newly arrived jobs. The reduction in prefilling time reflects the decoding time improvement observed in RE, as depicted in Figure 16.
Inference cost. We evaluate the resource cost based on the on-demand price of AWS EC2 instances [5, 6], i.e., \$5/hour per A100 GPU, \$0.0088/hour/GB for DRAM and \$0.000082/hour/GB for SSD. Figure 18 shows the total costs of different methods for completing the workload. Compared to RE, AS achieves significant cost savings for LLaMA-13B, LLaMA-65B, LLaMA-70B, and Falcon-40B, amounting to 70%, 43%, 66%, and 68%, respectively. These cost savings primarily stem from the reduced GPU time, as AS effectively reduces redundant prefilling for historical tokens and recomputation costs during context overflow, as depicted in Figure 16. AS employs cost-effective storage mediums including host memory and disks to cache the KV caches during inactive conversation sessions. The storage cost from the host memory and disks constitutes 16.4%, 9.0%, 9.0%, and 9.0% of the total cost in AS for LLaMA-13B, LLaMA-65B, LLaMA-70B, and Falcon-40B, respectively.
@@ 4.3.3 Scheduler-aware Fetching and Eviction
As shown in Figure 21(a), for the configuration of 128G/2T that indicates 128GB DRAM and 2TB SSD, AS outperforms LRU and FIFO in the overall cache hit rate by 27% and 31%, respectively. With the increased SSD capacity (128G/10T), AS achieves a remarkable hit rate of 86%, surpassing LRU (58%) and FIFO (48%). AS achieves high overall hit rates because AS is aware of the future KV cache access information to avoid evicting the KV caches that will be used in the future. The higher hit rates are translated to the reduced GPU time as shown in Figure 21(b), where AS achieves speedup up to 2.7$\times$. Analyzing the breakdown of hit rate, for the configuration of 128G/2T, LRU and FIFO only achieve 0.5% and 0.5% (too tiny to display) DRAM hit rates, with the remaining 12.4% and 9.0% being disk hit rates. Even with the overall hit rate increasing to 58% and 48% respectively for the larger capacity of 128G/10T, LRU and FIFO still exhibit limited DRAM hit rates of approximately 0.6% and 0.5% respectively. This is because LRU and FIFO lack awareness of future KV cache information and cannot pre-fetch KV caches from disks to host memory, thereby limiting their ability to improve DRAM hit rates. In contrast, AS achieves a cache hit rate of up to 86%, with over 99.6% of the hits occurring in DRAM due to its scheduler-aware policy.